show that the vertices represented by the complex numbers 6-i, 7+3i, 8+2i & 7-2i form a parallelogram
Answer by ikleyn(53763)


Let us introduce complex numbers
x = 6 - i;
y = 7 + 3i;
u = 8 + 2i;
v = 7 - 2i.

They are vertices of a quadrilateral in the complex plane.

Now,  calculate these differences of complex numbers:
y - z = 1 + 4i,     and
u - v = 1 + 4i.

The difference  y-z  represents the side  zy  of the quadrilateral  (the segment from the vertex  z  to the vertex  y).

The difference  u-v  represents the side  vu  of the quadrilateral  (the segment from the vertex  v  to the vertex  u).

As you see,  the complex numbers  y-z  and  u-v  are equal.
It means that the corresponding segments are parallel and have the same length.

Hence,  our quadrilateral is a parallelogram.

See my lessons on complex numbers in this site:
REVIEW of lessons on complex numbers.